OpenStudy (anonymous):

(-x^2) (-x^5) Please show work and how to do the problem. Thanks!

OpenStudy (anonymous):

a^m * a^n = a^(m+n)

OpenStudy (anonymous):

(-x)^2 + 5 = -x^7

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw